Mike has twice as many stamps as john has. After he gives john 6 stamps, he still has 8 more stamps than john does. How many stamps did mike have originally?
Found 3 solutions by addingup, ikleyn, MathTherapy:
Answer by addingup(3677)

M = 2J
M-6 = J+8 in this equation substitute for M:
2J-6 = J+8 subtract J and add 6 on both sides
J = 14 and the problem says that M has twice as many, so:
M = 14*2 = 28

Answer by ikleyn(52781)

.
M = 2J

M-6 = (J+6)+8 in this equation substitute for M:

2J-6 = J+14.       subtract J from both sides. Add 6 on both sides

J = 20 and the problem says that M has twice as many, so:

M = 20*2 = 40.


Correct answer is M = 40, J = 20.



Answer by MathTherapy(10552)


Mike has twice as many stamps as john has. After he gives john 6 stamps, he still has 8 more stamps than john does. How many stamps did mike have originally?
Correct answer: 
IGNORE anyone who says otherwise!